OpenStudy (mathhelp346):

Which is longer, the perimeter of a card table measuring 27 inches on a side or the circumference of a circular table measuring 33 inches in diameter?

OpenStudy (mathhelp346):

It's the card table right

OpenStudy (anonymous):

If the table is square (I have no familiarity with a 'card table,' sorry), then its perimeter is: \[ P=4s=4(27)=108\text{ in} \]Now, if we have a circle with diameter \(d\): \[ C=2\pi r=\pi d=\pi(33)\approx 103.67\text{ in} \]So, since \(P>C\), the card table has the longer of the two measurements.

OpenStudy (mathhelp346):

ty